ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 94, 1996

Akcoglu, M. and J. L. King An example of pointwise nonconvergence of iterated conditional expectation operators

179

Alon, N., B. Mohar and D. P. Sanders On acyclic colorings of graphs on surfaces

273

Bergen, J. and P. Grzeszczuk Invariants of Lie superalgebras acting on associative algebras

403

Cawley, E., B. Marcus and S. Tuncel Boundary measures of Markov chains

111

Cohen, J. M. and F. Colonna Extreme points of the Bloch space of a homogeneous tree

247

Colonna, F. See J. M. Cohen

Downey, R. and R. A. Shore Lattice embeddings below a nonlow_{2} recursively enumerable degree

221

Escauriaza, L. The L^{p} Dirichlet problem for small perturbations of the Laplacian

353

Ferrero, M. and E. R. Puczyłowski Prime ideals and radicals of centred extensions and tensor products

381

Galaktionov, V. A. and H. A. Levine On critical Fujita exponents for heat equations with nonlinear flux conditions on the boundary

125

Giménez, F. and V. Miquel On nonimmersibility of compact hypersurfaces into a ball of a simply connected space form

21

Grzeszczuk, P. See J. Bergen

Häggström, O. On phase transitions for subshifts of finite type  319

Jaworski, W. The asymptotic σ-algebra of a recurrent random walk on a locally compact group

201

Kenyon, R. and Y. Peres Hausdorff dimensions of sofic affineinvariant sets

157

King, J. L. See M. Akcoglu

Levin, G. and F. Przytycki External rays to periodic points  29

Levine, H. A. See V. A. Galaktionov

Manevitz, L. M. and S. Weinberger Discrete circle actions: A note using non-standard analysis

147

Marcus, B. See E. Cawley Miquel, V. See F. Giménez Mohar, B. See N. Alon Müller, P. Reducibility behavior of polynomials with varying

coefficients

59

Peres, Y. See R. Kenyon Power, S. C. On the Banach space isomorphism type of AF

C^{∗}-algebras and their triangular subalgebras

93

Przytycki, F. See G. Levin Puczyłowski, E. R. See M. Ferrero

Roberts, B. The theta correspondence for similitudes  285

Sanders, D. P. See N. Alon

Segal, D. A footnote on residually finite groups  1
Segal, D. Variations on polynomial subgroup growth  7

Shore, R. A. See R. Downey

Stellmacher, B. A characteristic subgroup of Σ_{4}-free groups  367

Tuncel, S. See E. Cawley Wang, L. Brauer–Manin obstruction to weak approximation on

abelian varieties

189

Weinberger, S. See L. M. Manevitz